the period of the function f(x)= |cos^3 3x| is pi/3. True or false with full explanation

1
Expert's answer
2022-03-15T18:27:18-0400

Solution


The function "y = \left| {\,{{\cos }^3}\left( {3x} \right)\,} \right|" is plotted and is shown below.





As per the plot above a couple of cycles has been plotted


It is obvious that the period is "p = \frac{\pi }{3}"


Hence the statement "the period of the function f(x)= |cos^3 3x| is pi/3", is TRUE